OpenStudy (anonymous):

7.Given the maximum value of the quadratic function f(x)=-2x^2-5x+p is 1/2,find the value of p.pls check whether my answer and working is right. @Directrix @Jhannybean @freckles @ganeshie8

OpenStudy (freckles):

\[-2x^2-5x+p \\ =-2(x^2+\frac{5}{2}x)+p \\ =-2(x^2+\frac{5}{2}x+(\frac{5}{2 \cdot 2})^2)+p+2(\frac{5}{2 \cdot 2})^2 \\ =-2(x+\frac{5}{2 \cdot 2})^2+p+2( \frac{5}{2 \cdot 2})^2\]

OpenStudy (freckles):

added in ((5/2)/2 )^2 to complete the square

OpenStudy (freckles):

but it was being multiply by -2 so -2(5/4)^2+2(5/4)^2 gives 0 so I actually added in zero

OpenStudy (freckles):

\[=-2(x+\frac{5}{4})^2+p+2 \frac{25}{16} \\-2(x+\frac{5}{4})^2+p+\frac{25}{8}\]

OpenStudy (freckles):

in general, \[y=ax^2+bx+c \\ y=a(x^2+\frac{b}{a}x)+c \\ y=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2 \\ y=a(x+\frac{b}{2a})^2+c-\frac{ab^2}{4a^2} \\ y=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}\] \[\frac{1}{2}=c-\frac{b^2}{4a} \\ \frac{1}{2}=p-\frac{(-5)^2}{4(-2)} \\ \frac{1}{2}=p-\frac{25}{-8} \\ \frac{1}{2}=p+\frac{25}{8}\] \[p=\frac{1}{2}-\frac{25}{8} \\ p=\frac{4}{8}-\frac{25}{8}\]

OpenStudy (freckles):

4-25 doesn't give -29

OpenStudy (freckles):

hmmm 4-25=-21 I don't think -21 and 8 have any common factors

OpenStudy (freckles):

do you agree the max value is p+25/8?

OpenStudy (freckles):

and also 1/2

OpenStudy (anonymous):

\[f(x)=-\left[ 2x^2+5x-p \right]\] \[=-\left[ (2x+\frac{ 5 }{ 2 })^2-\frac{ 25 }{ 4 }-p \right]\] \[=-2(x+\frac{ 5 }{ 4 })^2+\frac{ 25 }{ 4 }+p\] \[\frac{ 25 }{ 4 }+p=-\frac{ 1 }{ 2 }\] \[p =-\frac{ 25 }{ 4 }-\frac{ 1 }{ 2 }\] \[p=-\frac{ 27 }{ 4 }\]

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

Why do u need to multiply it by 2?

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[f(x)=-2x^2-5x+p \] factor out -2 from the first two terms only, don't touch the constant \(p\). you get \[f(x)=-2\left(x^2+\frac{5}{2}x\right)+p \] yes ?

ganeshie8 (ganeshie8):

which is same as \[f(x)=-2\left(x^2+2\cdot \frac{5}{4}x\right)+p \]

OpenStudy (anonymous):

\(\bf\huge\color{#ff0000}{T}\color{#ff2000}{h}\color{#ff4000}{a}\color{#ff5f00}{n}\color{#ff7f00}{k}~\color{#ffaa00}{Y}\color{#ffd400}{o}\color{#bfff00}{u}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @ganeshie8

OpenStudy (anonymous):

and @freckles